# Entanglement Entropy for Screened Interactions via Dimensional Mapping to Harmonic Oscillators

### v1

@@ II Map to coupled oscillator
$\frac{-\hbar^{2}}{2m_{1}}\frac{\partial^{2}\Psi(x_{1},x_{2})}{\partial x_{1}^{2}}-\frac{\hbar^{2}}{2m_{2}}\frac{\partial^{2}\Psi(x_{1},x_{2})}{\partial x_{2}^{2}}$
$-g^{2}\frac{e^{-\alpha m|x_{1}-x_{2}|}}{|x_{1}-x_{2}|}\Psi(x_{1},x_{2})=-B\Psi(x_{1},x_{2})$ (2)
@@
$x_{1}$ $=\frac{Mx_{c}+m_{2}x}{M}$ (4)
$x_{2}$ $=\frac{Mx_{c}-m_{1}x}{M}$ (5)
where $M=m_{1}+m_{2}$.
The schrodinger equation now becomes,
@@
$-\frac{\hbar^{2}}{2M}\frac{1}{g(x_{c})}\frac{\partial^{2}g(x_{c})}{\partial x_{c}^{2}}-\frac{\hbar^{2}}{2\mu}\frac{1}{f(x)}\frac{\partial^{2}f(x)}{\partial x_{c}^{2}}$
$-g^{2}\frac{e^{-\alpha m|x|}}{|x|}=-B$ (8)
@@
$\frac{\hbar^{2}}{2M}\frac{\partial^{2}g(x_{c})}{\partial x_{c}^{2}}=B_{c}g(x_{c})$ (9)
@@
While the first equation describes free particle motion, our interest lies in the second equation which can be used for describing both the bound as well as scattering states. Since $x$ and $x_{c}$ both are now independent coordinates, we can freely use non-linear transformation without the loss of generality and restore the original coordinates at the end of derivation to obtain the entangled wavefunction and density operator.
@@
$-\frac{\hbar^{2}}{2\mu}\left(\frac{d^{2}}{dr^{2}}-\frac{(d-1)(d-3)}{4r^{2}}-\frac{l(l+1)}{r^{2}}\right)u(r)$
$+\frac{1}{2}\mu\omega^{2}r^{2}u(r)=Eu(r)$ (13)
@@
$\frac{-\hbar^{2}}{2\mu}\left(-\frac{\phi}{4\rho^{3/2}}+\sqrt{\rho}\phi^{\prime\prime}-\frac{\phi}{2\rho^{3/2}}\right)$ $+4B_{r}\rho^{2}\sqrt{\rho}\phi$
$=4g^{2}e^{-\alpha m\rho^{2}}\sqrt{\rho}\phi$ (17)
@@
This equation is similar to 4-d radial harmonic oscillator equation Eq.(II) for $l=0$, except for the exponential term on R.H.S which also prevents it from being an eigenvalue equation. This can be resolved by considering perturbations to the 4-d harmonic oscillator by a series of anharmonic terms. We want to find the perturbed ground state wavefunction which would allow us to compute the density operator and then the entanglement entropy. In the rest of this work, we will study the effects of anharmonic perturbations to the entanglement entropy. A typical perturbation term in the series is of the form $\alpha^{n}m^{n}\rho^{2n}/n!$.
III Perturbed Ground State
@@
$\frac{-\hbar^{2}}{2\mu}\left(\frac{d^{2}\phi}{d\rho^{2}}-\frac{3}{4}\frac{\phi}{\rho^{2}}\right)+4B_{r}\rho^{2}\phi\,$
$=4g^{2}\phi-4g^{2}\alpha m\rho^{2}\phi+\frac{1}{2!}4g^{2}\alpha^{2}m^{2}\rho^{4}\phi$ (18)
@@
$\frac{-\hbar^{2}}{2\mu}\left(\frac{d^{2}\phi}{d\rho^{2}}-\frac{3}{4}\frac{\phi}{\rho^{2}}\right)+4B_{r}^{\prime}\rho^{2}\phi-2g^{2}\alpha^{2}m^{2}\rho^{4}\phi$
$=4g^{2}\phi$ (19)
@@
$\omega_{0}\equiv\sqrt{\frac{8B_{r}^{\prime}}{\mu}},\qquad\beta\equiv\frac{\mu\omega_{0}}{\hbar}.$ (21)
@@
where $n_{r}$ is the radial quantum number, $\ell$ is the angular momentum, and $L_{n_{r}}^{(\ell+1)}$ are associated Laguerre polynomials. Since the mapping leads to a spherically symmetric effective problem, we restrict attention to the $\ell=0$ sector throughout.
The first-order correction to the $n$th eigenstate due to a perturbation $H^{\prime}$ is given by standard Rayleigh–Schrödinger perturbation theory,
@@
Because $\rho^{4}$ is a scalar operator, it preserves angular momentum and couples only states with the same $\ell$. Consequently, only $\ell=0$ radial excitations contribute to the correction of the ground state.
Introducing the dimensionless variable $x=\beta\rho^{2}$, the relevant matrix elements reduce to
@@
implying that only the $n_{r}=1,2$ states contribute at this order.
The corresponding energy denominators are
@@
$a_{1}$ $=-\frac{12\sqrt{2}\,g^{2}\pi^{2}\alpha^{2}m^{2}}{\beta^{2}\hbar\omega_{0}},$ (31)
$a_{2}$ $=\frac{2\sqrt{3}\,g^{2}\pi^{2}\alpha^{2}m^{2}}{\beta^{2}\hbar\omega_{0}}.$ (32)
@@
$a_{1}$ $=-12\sqrt{2}\,\frac{\hbar g^{2}\pi^{2}\alpha^{2}m^{2}}{\mu^{2}\omega_{0}^{3}},$ (33)
$a_{2}$ $=2\sqrt{3}\,\frac{\hbar g^{2}\pi^{2}\alpha^{2}m^{2}}{\mu^{2}\omega_{0}^{3}}.$ (34)
@@
$c_{0}$ $=\sqrt{2}\,\beta-18\sqrt{2}\,\frac{g^{2}\pi^{2}\alpha^{2}m^{2}}{\hbar\omega_{0}\beta},$ (37)
$c_{2}$ $=6\sqrt{2}\,\frac{g^{2}\pi^{2}\alpha^{2}m^{2}}{\hbar\omega_{0}},$ (38)
$c_{4}$ $=\sqrt{2}\,\beta\,\frac{g^{2}\pi^{2}\alpha^{2}m^{2}}{\hbar\omega_{0}}.$ (39)
This explicit polynomial–Gaussian structure of the corrected wavefunction will be central to the construction of the reduced density matrix and the subsequent computation of entanglement entropy.
[…]
$\phi_{0}^{1}(r_{1},r_{2})=$ $e^{-\frac{1}{2}\beta(r_{1}-r_{2})^{2}}\Biggl[\sqrt{2}\beta-18\sqrt{2}\,\frac{g^{2}\pi^{2}\alpha^{2}m^{2}}{\hbar\omega_{0}\beta}$
$\qquad+6\sqrt{2}\frac{g^{2}\pi^{2}\alpha^{2}m^{2}}{\hbar\omega_{0}}(r_{1}-r_{2})^{2}$
$\qquad+2\sqrt{2}\beta\frac{g^{2}\pi^{2}\alpha^{2}m^{2}}{2\hbar\omega_{0}}(r_{1}-r_{2})^{4}\Biggr]$ (40)
@@
IV Reduced Density Matrix
@@
$\rho(s)=2\pi^{2}e^{-\beta s^{2}/4}\int_{0}^{\infty}y^{3}dye^{-\beta y^{2}}\bigg[\pi c_{0}^{2}/2+$ (45)
$\pi c_{0}c_{2}(y^{2}+s^{2}/4)+\pi c_{0}c_{4}[(y^{2}+s^{2}/4)^{2}+y^{2}s^{2}/4]\bigg]$
@@
$\rho(s)$ $=c_{0}\pi^{3}e^{-\beta s^{2}/4}$
$\times\frac{\left(96c_{4}+\beta\left(8\beta c_{0}+32c_{2}+4(\beta c_{2}+6c_{4})s^{2}+\beta c_{4}s^{4}\right)\right)}{16\beta^{4}}$
@@
$\tilde{\rho}(k)=$ $\frac{e^{-k^{2}/\beta}\pi^{7/2}}{\beta^{9/2}\hbar\omega_{0}}\Big(2\,\beta^{4}\,\hbar\,\omega_{0}+2\,g^{2}m^{2}\pi^{2}\alpha^{2}$
$\times\big[27\beta^{2}-60\beta k^{2}+4k^{4}\big]\Big)$ (50)
Since the reduced density matrix is translationally invariant and has a continuous spectrum, we will define the entanglement entropy later by using normalized momentum-space eigenvalue density given by
$\rho(k)=$ $\frac{e^{-k^{2}/\beta}}{2\beta^{9/2}\pi^{1/2}\hbar\omega_{0}}\Big(2\,\beta^{4}\,\hbar\,\omega_{0}+2\,g^{2}m^{2}\pi^{2}\alpha^{2}$
$\times\big[27\beta^{2}-60\beta k^{2}+4k^{4}\big]\Big)$ (51)
@@
$\beta_{0}$ $=\frac{\mu}{\hbar}\sqrt{\frac{8B_{r}}{\mu}},$ (54)
$\beta_{1}$ $=\frac{g^{2}m}{2B_{r}}\,\beta_{0},$ (55)
$\beta_{2}$ $=-\frac{1}{8}\left(\frac{g^{2}m}{B_{r}}\right)^{2}\beta_{0}.$ (56)
@@
$e^{-k^{2}/\beta(\alpha)}$ $=e^{-k^{2}/\beta_{0}}\Big[1+\alpha\,\frac{\beta_{1}k^{2}}{\beta_{0}^{2}}+\alpha^{2}\!\left(\frac{\beta_{2}k^{2}}{\beta_{0}^{2}}-\frac{\beta_{1}^{2}k^{2}}{\beta_{0}^{3}}\right$. (57)
$\qquad\qquad\left.+\frac{\beta_{1}^{2}k^{4}}{2\beta_{0}^{4}}\right)+\mathcal{O}(\alpha^{3})\Big]$
@@
Only after all $\beta$-dependent terms have been expanded and collected can the eigenvalue density be written in the schematic form
@@
It is important to note the consequences of the normalization of the eigenvalue density. By construction, the reduced eigenvalue $\rho(k)$ is normalized to unity,
@@
$\int dk\,\rho^{(0)}(k)$ $=1,$ (60)
$\int dk\,\rho^{(1)}(k)$ $=0,$ (61)
$\int dk\,\rho^{(2)}(k)$ $=0,$ (62)
and similarly for all higher orders. The zeroth-order condition Eq. (60) fixes the normalization of the unperturbed Gaussian eigenvalue, while the higher-order constraints Eqs. (61)–(62) ensure that perturbative corrections only redistribute spectral weight without changing the total probability.
These relations play a crucial role in the entropy calculation. In particular, they guarantee that linear and quadratic corrections to the von Neumann entropy may be consistently simplified using integration by parts and Gaussian moment identities, and they ensure that spurious constant contributions do not appear in $S^{(1)}$ and $S^{(2)}$. Throughout the remainder of this work, we explicitly enforce the normalization conditions Eq. (61) and Eq. (62), which serve as nontrivial consistency checks on the perturbative density matrix and its Gaussian and non-Gaussian components.
As in Eq.(IV), the eigenvalue of the reduced density matrix obtained from the partially traced ground state takes the form
@@
This correction originates entirely from the harmonic ($\rho^{2}$) perturbation and corresponds to a renormalization of the Gaussian width. It does not introduce genuinely non-Gaussian correlations.
At order $\alpha^{2}$, two distinct contributions arise.
@@
$A_{0}$ $=\frac{3}{8}\frac{\beta_{1}^{2}}{\beta_{0}^{2}}-\frac{1}{2}\frac{\beta_{2}}{\beta_{0}},$ (68)
$A_{2}$ $=\frac{\beta_{2}}{\beta_{0}^{2}}-\frac{3}{2}\frac{\beta_{1}^{2}}{\beta_{0}^{3}},$ (69)
$A_{4}$ $=\frac{1}{2}\frac{\beta_{1}^{2}}{\beta_{0}^{4}}$ (70)
@@
V Perturbative evaluation of entanglement entropy

### v2

@@ 2 Map to coupled oscillator
$\frac{-\hbar^{2}}{2m_{1}}\frac{\partial^{2}\Psi(x_{1},x_{2})}{\partial x_{1}^{2}}-\frac{\hbar^{2}}{2m_{2}}\frac{\partial^{2}\Psi(x_{1},x_{2})}{\partial x_{2}^{2}}-g^{2}\frac{e^{-\alpha m|x_{1}-x_{2}|}}{|x_{1}-x_{2}|}\Psi(x_{1},x_{2})=-B\Psi(x_{1},x_{2})$ (2)
@@
$x_{1}=\frac{Mx_{c}+m_{2}x}{M},\quad x_{2}=\frac{Mx_{c}-m_{1}x}{M}$ (4)
where $M=m_{1}+m_{2}$. The schrodinger equation now becomes,
@@
$-\frac{\hbar^{2}}{2M}\frac{1}{g(x_{c})}\frac{\partial^{2}g(x_{c})}{\partial x_{c}^{2}}-\frac{\hbar^{2}}{2\mu}\frac{1}{f(x)}\frac{\partial^{2}f(x)}{\partial x_{c}^{2}}-g^{2}\frac{e^{-\alpha m|x|}}{|x|}=-B$ (7)
@@
$\frac{\hbar^{2}}{2M}\frac{\partial^{2}g(x_{c})}{\partial x_{c}^{2}}=E_{c}\,g(x_{c})$ (8)
@@
after separation, the total energy is written as $E=E_{c}-B_{r}$, where $E_{c}$ being the center of mass kinetic energy and $B_{r}>0$ is the binding energy associated with the relative motion. While the first equation describes free particle motion, our interest lies in the second equation which can be used for describing both the bound as well as scattering states. Since $x$ and $x_{c}$ both are now independent coordinates, we can freely use non-linear transformation without the loss of generality and restore the original coordinates at the end of derivation to obtain the entangled wavefunction and density operator.
@@
$-\frac{\hbar^{2}}{2\mu}\left(\frac{d^{2}}{dr^{2}}-\frac{(d-1)(d-3)}{4r^{2}}-\frac{l(l+1)}{r^{2}}\right)u(r)+\frac{1}{2}\mu\omega^{2}r^{2}u(r)=Eu(r)$ (12)
@@
$\frac{-\hbar^{2}}{2\mu}\left(-\frac{\phi}{4\rho^{3/2}}+\sqrt{\rho}\phi^{\prime\prime}-\frac{\phi}{2\rho^{3/2}}\right)+4B_{r}\rho^{2}\sqrt{\rho}\phi=4g^{2}e^{-\alpha m\rho^{2}}\sqrt{\rho}\phi$ (16)
@@
This equation is similar to 4-d radial harmonic oscillator equation eq.(12) for $l=0$, except for the exponential term which also prevents it from being an eigenvalue equation. This can be resolved by considering perturbations to the 4-d harmonic oscillator by a series of anharmonic terms. We want to find the perturbed ground state wavefunction which would allow us to compute the density operator and then the entanglement entropy. In the rest of this work, we will study the effects of anharmonic perturbations to the entanglement entropy. A typical perturbation term in the series is of the form $\alpha^{n}m^{n}\rho^{2n}/n!$.
3 Perturbed Ground State
@@
$\frac{-\hbar^{2}}{2\mu}\left(\frac{d^{2}\phi}{d\rho^{2}}-\frac{3}{4}\frac{\phi}{\rho^{2}}\right)+4B_{r}\rho^{2}\phi\,=4g^{2}\phi-4g^{2}\alpha m\rho^{2}\phi+\frac{1}{2!}4g^{2}\alpha^{2}m^{2}\rho^{4}\phi$ (17)
@@
$\frac{-\hbar^{2}}{2\mu}\left(\frac{d^{2}\phi}{d\rho^{2}}-\frac{3}{4}\frac{\phi}{\rho^{2}}\right)+4B_{r}^{\prime}\rho^{2}\phi-2g^{2}\alpha^{2}m^{2}\rho^{4}\phi=4g^{2}\phi$ (18)
@@
$\omega_{0}\equiv\sqrt{\frac{8B_{r}^{\prime}}{\mu}},\quad\beta\equiv\frac{\mu\omega_{0}}{\hbar}.$ (20)
@@
where $n_{r}$ is the radial quantum number, $\ell$ is the angular momentum, and $L_{n_{r}}^{(\ell+1)}$ are associated Laguerre polynomials. Since the mapping leads to a spherically symmetric effective problem, we restrict attention to the $\ell=0$ sector throughout. The first-order correction to the $n$th eigenstate due to a perturbation $H^{\prime}$ is given by standard Rayleigh–Schrödinger perturbation theory,
@@
Because $\rho^{4}$ is a scalar operator, it preserves angular momentum and couples only states with the same $\ell$. Consequently, only $\ell=0$ radial excitations contribute to the correction of the ground state. Introducing the dimensionless variable $x=\beta\rho^{2}$, the relevant matrix elements reduce to
@@
implying that only the $n_{r}=1,2$ states contribute at this order. The corresponding energy denominators are
@@
$a_{1}=-\frac{12\sqrt{2}\,g^{2}\pi^{2}\alpha^{2}m^{2}}{\beta^{2}\hbar\omega_{0}},\quad a_{2}=\frac{2\sqrt{3}\,g^{2}\pi^{2}\alpha^{2}m^{2}}{\beta^{2}\hbar\omega_{0}}.$ (30)
@@
$a_{1}=-12\sqrt{2}\,\frac{\hbar g^{2}\pi^{2}\alpha^{2}m^{2}}{\mu^{2}\omega_{0}^{3}},\quad a_{2}=2\sqrt{3}\,\frac{\hbar g^{2}\pi^{2}\alpha^{2}m^{2}}{\mu^{2}\omega_{0}^{3}}.$ (31)
@@
$c_{0}=\sqrt{2}\,\beta-18\sqrt{2}\,\frac{g^{2}\pi^{2}\alpha^{2}m^{2}}{\hbar\omega_{0}\beta},\,c_{2}=6\sqrt{2}\,\frac{g^{2}\pi^{2}\alpha^{2}m^{2}}{\hbar\omega_{0}},\,c_{4}=\sqrt{2}\,\beta\,\frac{g^{2}\pi^{2}\alpha^{2}m^{2}}{\hbar\omega_{0}}.$ (34)
This explicit polynomial–Gaussian structure of the corrected wavefunction will be central to the construction of the reduced density matrix and the subsequent computation of entanglement entropy. So the corrected wavefunction in terms of radial coordinates $r_{1}$ and $r_{2}$ is given by
$\phi_{0}^{1}(r_{1},r_{2})$ $=\,e^{-\frac{1}{2}\beta(r_{1}-r_{2})^{2}}\Biggl[\sqrt{2}\beta-18\sqrt{2}\,\frac{g^{2}\pi^{2}\alpha^{2}m^{2}}{\hbar\omega_{0}\beta}+6\sqrt{2}\frac{g^{2}\pi^{2}\alpha^{2}m^{2}}{\hbar\omega_{0}}(r_{1}-r_{2})^{2}$
$\quad+2\sqrt{2}\beta\frac{g^{2}\pi^{2}\alpha^{2}m^{2}}{2\hbar\omega_{0}}(r_{1}-r_{2})^{4}\Biggr]$ (35)
@@
4 Reduced Density Matrix
@@
$\rho(s)=2\pi^{2}e^{-\beta s^{2}/4}\int_{0}^{\infty}y^{3}dye^{-\beta y^{2}}$ $\bigg[\pi c_{0}^{2}/2+\pi c_{0}c_{2}(y^{2}+s^{2}/4)+\pi c_{0}c_{4}[(y^{2}+s^{2}/4)^{2}$ (40)
$+y^{2}s^{2}/4]\bigg]$
@@
$\rho(s)=c_{0}\pi^{3}e^{-\beta s^{2}/4}\frac{\left(96c_{4}+\beta\left(8\beta c_{0}+32c_{2}+4(\beta c_{2}+6c_{4})s^{2}+\beta c_{4}s^{4}\right)\right)}{16\beta^{4}}$ (41)
@@
$\tilde{\rho}(k)=\frac{e^{-k^{2}/\beta}\pi^{7/2}}{\beta^{9/2}\hbar\omega_{0}}\Big(2\,\beta^{4}\,\hbar\,\omega_{0}+2\,g^{2}m^{2}\pi^{2}\alpha^{2}\big[27\beta^{2}-60\beta k^{2}+4k^{4}\big]\Big)$ (45)
Since the reduced density matrix is translationally invariant and has a continuous spectrum, we will define the entanglement entropy later by using normalized momentum-space eigenvalue given by
$\rho(k)=\frac{e^{-k^{2}/\beta}}{2\beta^{9/2}\pi^{1/2}\hbar\omega_{0}}\Big(2\,\beta^{4}\,\hbar\,\omega_{0}+2\,g^{2}m^{2}\pi^{2}\alpha^{2}\big[27\beta^{2}-60\beta k^{2}+4k^{4}\big]\Big)$ (46)
@@
$\beta_{0}=\frac{\mu}{\hbar}\sqrt{\frac{8B_{r}}{\mu}},\quad\beta_{1}=\frac{g^{2}m}{2B_{r}}\,\beta_{0},\quad\beta_{2}=-\frac{1}{8}\left(\frac{g^{2}m}{B_{r}}\right)^{2}\beta_{0}.$ (49)
@@
$e^{-k^{2}/\beta(\alpha)}$ $=e^{-k^{2}/\beta_{0}}\left[1+\alpha\,\frac{\beta_{1}k^{2}}{\beta_{0}^{2}}+\alpha^{2}\!\left(\frac{\beta_{2}k^{2}}{\beta_{0}^{2}}-\frac{\beta_{1}^{2}k^{2}}{\beta_{0}^{3}}+\frac{\beta_{1}^{2}k^{4}}{2\beta_{0}^{4}}\right)+\mathcal{O}(\alpha^{3})\right]$ (50)
@@
Only after all $\beta$-dependent terms have been expanded and collected can the eigenvalue be written in the schematic form
@@
It is important to note the consequences of the normalization of the momentum space eigenvalue. By construction, the reduced eigenvalue $\rho(k)$ is normalized to unity,
@@
$\int dk\,\rho^{(0)}(k)=1,\quad\int dk\,\rho^{(1)}(k)=0,\quad\int dk\,\rho^{(2)}(k)=0,$ (53)
and similarly for all higher orders. The zeroth-order condition in eq. (53) fixes the normalization of the unperturbed Gaussian eigenvalue, while the higher-order constraints in eq.(53) ensure that perturbative corrections only redistribute spectral weight without changing the total probability.
These relations play a crucial role in the entropy calculation. In particular, they guarantee that linear and quadratic corrections to the von Neumann entropy may be consistently simplified using integration by parts and Gaussian moment identities, and they ensure that spurious constant contributions do not appear in $S^{(1)}$ and $S^{(2)}$. Throughout the remainder of this work, we explicitly enforce the normalization conditions in eq.(53), which serve as nontrivial consistency checks on the perturbative density matrix and its Gaussian and non-Gaussian components.
As in eq.(46), the eigenvalue of the reduced density matrix obtained from the partially traced ground state takes the form
@@
This correction originates entirely from the harmonic ($\rho^{2}$) perturbation and corresponds to a renormalization of the Gaussian width. It does not introduce genuinely non-Gaussian correlations. At order $\alpha^{2}$, two distinct contributions arise.
@@
$A_{0}=\frac{3}{8}\frac{\beta_{1}^{2}}{\beta_{0}^{2}}-\frac{1}{2}\frac{\beta_{2}}{\beta_{0}},\quad A_{2}=\frac{\beta_{2}}{\beta_{0}^{2}}-\frac{3}{2}\frac{\beta_{1}^{2}}{\beta_{0}^{3}},\quad A_{4}=\frac{1}{2}\frac{\beta_{1}^{2}}{\beta_{0}^{4}}$ (59)
@@
5 Perturbative evaluation of entanglement entropy